OpenStudy (anonymous):

Find the equation of the planes through the straight line 2x-y+3z+2=0=3x+2y-z+3 parallel to the co-ordinate axes.

OpenStudy (amistre64):

2x-y+3z+2 = 0 = 3x+2y-z+3 , sounds like the line is the intersection of two planes what is a coordinate axes?

OpenStudy (anonymous):

x,y and z axes

OpenStudy (amistre64):

so three different planes?

OpenStudy (amistre64):

define 2 points on the line of intersection; let x=0, and solve the zy parts, then let z = 0 and solve the xy parts .... or whatever combination of 0 this one and solve the other 2 that fits your fancy 2x-y+3z+2 = 0 3x+2y-z+3 = 0 with 2 points, we can define a line between them

OpenStudy (anonymous):